# Untangling Hate Speech Definitions: A Semantic Componential Analysis Across Cultures and Domains

### v1

@@ Abstract
Hate speech relies heavily on cultural influences, leading to varying individual interpretations. For that reason, we propose a Semantic Componential Analysis (SCA) framework for a cross-cultural and cross-domain analysis of hate speech definitions. We create the first dataset of definitions derived from five domains: online dictionaries, research papers, Wikipedia articles, legislation, and online platforms, which are later analyzed into semantic components. Our analysis reveals that the components differ from definition to definition, yet many domains borrow definitions from one another without taking into account the target culture. We conduct zero-shot model experiments using our proposed dataset, employing three popular open-sourced LLMs to understand the impact of different definitions on hate speech detection. Our findings indicate that LLMs are sensitive to definitions: responses for hate speech detection change according to the complexity of definitions used in the prompt.
@@ 1 Introduction
Inspired by the compositionality principle Hinzen et al. (2012), we introduce a component-annotated resource for hate speech definitions, the HateDefCon dataset. Figure 1 shows our HateDefCon creation pipeline. We propose a Semantic Componential Analysis (SCA), in which we define a hate speech definitional component as a fundamental element or criterion used to define what constitutes hate speech in terms of target, intention/purpose, and act/means. We define definitional hate speech domains as the contexts where hate speech definitions emerge. This study focuses on five such domains: legislation, Wikipedia, online dictionaries, research papers, and conduct policies from online platforms or technological companies. In addition, we look at the cultural representation of hate speech. We use culture, as the term that encompasses language, ideas, beliefs, customs, codes, institutions, tools, techniques, among other elements. [^1] https://www.britannica.com/topic/culture We highlight the need for cross-cultural and cross-domain approaches to define hate speech and argue that such definitions should be context-specific, be it cultural, legal, or academic. We advocate for grounding hate speech definitions within these particular domains. This is in line with best practices for tackling subjective tasks Rottger et al. (2022), in which the guideline - or the definition - chosen should consider the downstream use, i.e. the context.
In this work, we focus on three research questions: (1) What are the differences among various definitions of hate speech? (2) What is the diversity of these definitions? (3) How do definitions with different components affect LLM predictions? Our contributions are both theoretical and practical. On the theoretical front, our cross-cultural and cross-domain analysis of definitions shows significant variation in components, ranging from broad definitions to highly specific ones. Even among the more detailed definitions, which address aspects like the target of hate speech, the intent, and the methods of expressing it, there are differences in their components. On the practical side, we assess whether LLMs respond better to certain definitions, potentially revealing underlying biases. Our results reveal that definitions vary in their components while domains also borrow definitions from one another. When research is culturally specific, borrowing from other domains can be problematic, as it may introduce elements that do not align with the intended cultural context.
@@
  • We propose a semantically informed framework and use it to create HateDefCon, the first resource for hate speech definitions.
@@ 4 The HateDefCon Dataset
We analyze and compare hate speech definitions via SCA to identify potential cross-domain and cross-cultural differences. We collect definitions from five different domains in which we can find hate speech definitions (§4.1). We provide information about the respective sources and a final corpus analysis to gain cross-domain and cross-cultural insights (§4.2). Figure 1 summarizes the HateDefCon dataset creation pipeline. We collect definitions from the selected domains, we extract keywords which we use to annotate as components for each definition.
@@ 4.1 Data
We source definitions from the Global Handbook of Hate Speech Laws website, a comprehensive resource that provides access to existing hate speech legislation from countries worldwide, including the United Nations and European Union levels. [^2] https://futurefreespeech.org/global-handbook-on-hate-speech-laws/. The legislations are already available in English. We exclude countries for which their legislative texts were not available. The full list of the countries is available in Appendix B.
@@
We use the Hate Speech Dataset Catalogue to navigate through research articles and datasets, focusing specifically on definitions of hate speech. [^4] https://hatespeechdata.com/ We do not include other related terms like toxicity or abusive language, as our primary focus is hate speech.
@@ 4.2 Component Annotation
To ensure objectivity in the annotation process and a very fine-grained representation of the components, we establish that a definition contains a given component if a derivative of the component’s word appears in the annotation. For instance, if the word ‘abusive’ is present in the definition, the annotator marks the component ‘abuse’ as present. A challenge we encounter during annotation is the treatment of synonyms. For instance, the words ‘harm’ and ‘hurt’ can be considered synonymous, but they are not derivatives of the same root word. To maintain consistency and avoid subjective interpretation, which could lead to different grouping of the synonyms by the annotators, we only annotate derivatives of the target term (e.g., marking ‘harm’ only if words like ‘harmful’ or ‘harmed’ appear). Additionally, missing components are labeled as ‘undefined component’. This process allows us to comprehensively gather all components deemed crucial by the annotators.
To finalize the dataset, we keep all annotated components, as a manual inspection showed that most of the disagreement derived from the fact that one of the annotators missed a component. The average inter-annotator agreement (IAA) measured via Cohen’s kappa McHugh (2012) is 0.64 (see Appendix D for more details). Figure 2 reports the component hierarchy resulting from the annotation study (see Appendix E for a fine-grained version of the hierarchy).
@@ 5 Case Study: Definition Comparison
Domain No. Instances Culture Distribution Components
@@
Table 1: Number of instances, culture distribution, and top 5 components per domain.
@@
We employ SCA to describe potential inconsistencies between collected hate speech definitions and the cultural reality through a real case study. Mulki et al. (2019) develop a Levantine Hate Speech and Abusive Twitter dataset in an attempt to bring into the spotlight less-spoken Arabic varieties. However, in their paper, the authors refer to the hate speech definition by Nockleby (2000), which defines hate speech as “any communication that disparages a person or a group on the basis of some characteristic such as race, color, ethnicity, gender, sexual orientation, nationality, religion, or other characteristic”. Levantine Arabic is spoken in many Middle Eastern countries, such as Syria, Jordan, and Israel. Our SCA framework allows us to see if the definitions available for these countries overlap with the one used in the Mulki et al. (2019) paper (Table 2). While they all differ in the intent and act, some targets overlap, although the definition from Mulki et al. (2019) is more comprehensive, considering also sexual orientation and gender, which are not taken into account by the other definitions. Israel’s legislation seems to have the most detailed provisions, including concepts of hostility and cultural control, whereas Syria’s legislation is narrowly focused on race, and Jordan’s legislation emphasizes physical action and terrorism. This variation can affect annotation and prompting, resulting in varied interpretations of hate-related content. Therefore, tailoring prompts to align with specific regional definitions is essential for achieving consistent model behavior for a specific culture.
@@ 6 LLM Sensitivity to Definitions
We experiment with three open-source popular state-of-the-art LLMs: Llama3666Llama3-8B-instruct., Mistral2777Mistral-7B-Instruct-v0.2., and Phi3-mini. [^8] Phi-3-mini-4k-instruct. We consider the Gab Hate Corpus (GHC) on online hate speech conversations Kennedy et al. (2022). GHC consists of 27,665 posts from the social network service gab.com and is annotated for the presence of “hate-based rhetoric” by a minimum of three annotators. The posts are annotated based on a coding typology created by synthesizing definitions of hate speech from legal precedents, existing hate speech coding frameworks, and insights from psychology and sociology. This typology includes hierarchical labels that denote dehumanizing and violent speech, as well as indicators related to targeted groups and rhetorical framing. This dataset captures various dimensions of hate speech, making it well-suited for a detailed and fine-grained analysis and testing with multiple definitions. Although GHC is annotated in a multi-class classification, we consider a binary setting: hate vs not hate. To limit computational overhead, we select 500 instances from the corpus, equally divided between the two classes.
@@
We perform our experiments in a zero-shot setting to understand the impact of different definitions (i.e. different components) on model performance without any additional fine-tuning or prompting strategies. To prompt the models, we use three definitions of hate speech. We selected the definitions deliberately after manual inspection, aiming to assess whether the varying degrees of component coverage influence the prompting outcomes. This approach includes one definition with medium component coverage, a highly detailed one, and a very general one. We use three definitions: Dghc comes from the GHC dataset, it includes several components but is not the most comprehensive one; Dwiki is the definition provided by the Macedonian Wikipedia page, it offers a more detailed and varied perspective; and Ddict is a general definition from the Merriam-Webster dictionary (see Appendix F for more details). We set the temperature to zero to exclude variations in the generated response.
@@ 7 Results
Table 3 summarizes the results of the classification task according to the three definitions. Although Llama3 performs consistently better than other models, the complexity of the definitions affects the outputs of the three models differently. In Llama3, the comprehensiveness of the definition is directly proportional to the performance: the more comprehensive the definition is, the more the performance increases; in Mistral2, we observe a reverse tendency. Phi3-mini outperforms Mistral2 in all settings, although being half the size of Mistral2. Additionally, in Phi3-mini, we observe the smallest variance of responses based on the definitions, with Dghc performing the best.
There are cases in which Llama3 refuses to answer for safeguarding the generation process over harmful content. This tendency is also proportional to the completeness of the definition. The less comprehensive the definition, the more the model refuses to answer it. For Dwiki, the model refuses to answer to 1% of times, 1.6% for Dghc, and 5% for Ddict.
@@
Table 3: Classification performance using difference hate speech definitions.
@@
We extract all instances for which the answer of Llama3 varies according to the definition, regardless of the ground truth. Dghc and Dwiki differ in 54 instances, while Ddict differs from Dghc and Dwiki in 48 instances each. We observe that when the model is prompted with Dwiki, it tends to identify more personal attacks, which are often discarded in Dghc and Ddict as false negatives since such definitions identify group of people as targets of HS, rather than individuals. Moreover, with Dwiki, LLMs tend to identify more instances that are hateful with respect to political views and undocumented migrants. Indeed, the majority of instances that are correctly identified by Dwiki and misclassified by Dghc and Ddict, contain terms like liberals, communists, and illegal aliens with a negative connotation.
@@ Appendix E Complete Component Hierarchy and Further Statistics
Figure 4: Top 20 most frequent components across all definitions.

### v2

@@ Abstract
Hate speech relies heavily on cultural influences, leading to varying individual interpretations. For that reason, we propose a Semantic Componential Analysis (SCA) framework for a cross-cultural and cross-domain analysis of hate speech definitions. We create the first dataset of hate speech definitions encompassing 493 definitions from more than 100 cultures, drawn from five key domains: online dictionaries, academic research, Wikipedia, legal texts, and online platforms. By decomposing these definitions into semantic components, our analysis reveals significant variation across definitions, yet many domains borrow definitions from one another without taking into account the target culture. We conduct zero-shot model experiments using our proposed dataset, employing three popular open-sourced LLMs to understand the impact of different definitions on hate speech detection. Our findings indicate that LLMs are sensitive to definitions: responses for hate speech detection change according to the complexity of definitions used in the prompt.
@@ 1 Introduction
Inspired by the compositionality principle Hinzen et al. (2012), we introduce a component-annotated resource for hate speech definitions, the HateDefCon dataset. Figure 1 shows our HateDefCon creation pipeline. We propose a Semantic Componential Analysis (SCA), in which we define a hate speech definitional component as a fundamental element or criterion used to define what constitutes hate speech in terms of target, intention/purpose, and act/means. We define definitional hate speech domains as the contexts where hate speech definitions emerge. This study focuses on five such domains: legislation, Wikipedia, online dictionaries, research papers, and conduct policies from online platforms or technological companies. In addition, we look at the cultural representation of hate speech. We use culture, as the term that encompasses language, ideas, beliefs, customs, codes, institutions, tools, techniques, among other elements. [^1] https://www.britannica.com/topic/culture Distinguishing between culture and language in hate speech definitions is challenging because languages are not confined to a single culture or country. For example, Arabic is spoken across multiple nations with diverse cultural norms, making it difficult to attribute a hate speech definition to a specific cultural context. Additionally, if definitions are collected in a particular language without information on the cultural or national background, it becomes harder to account for variations in meaning, intent, and societal impact across different regions. Unfortunately, most definitions lack any kind of cultural information. We highlight the need for cross-cultural and cross-domain approaches to define hate speech and argue that such definitions should be context-specific, be it cultural, legal, or academic. We advocate for grounding hate speech definitions within these particular domains. This is in line with best practices for tackling subjective tasks Rottger et al. (2022), in which the guideline - or the definition - chosen should consider the downstream use, i.e. the context.
In this work, we focus on three research questions: (1) What are the differences among various definitions of hate speech? (2) What is the diversity of these definitions? (3) How do definitions with different components affect the predictions of Large Language Models (LLMs)? Our contributions are both theoretical and practical. On the theoretical front, our cross-cultural and cross-domain analysis of definitions shows significant variation in components, ranging from broad definitions to highly specific ones. Even among the more detailed definitions, which address aspects like the target of hate speech, the intent, and the methods of expressing it, there are differences in their components. On the practical side, we assess whether LLMs respond better to certain definitions, potentially revealing underlying biases. Our results reveal that definitions vary in their components while domains also borrow definitions from one another. When research is culturally specific, borrowing from other domains can be problematic, as it may introduce elements that do not align with the intended cultural context.
@@
  • We propose a semantically informed framework and use it to create HateDefCon, the first resource for hate speech definitions that amounts to 493 items.
@@ 4 The HateDefCon Dataset
We analyze and compare hate speech definitions via SCA to identify potential cross-domain and cross-cultural differences. We collect 493 definitions from five different domains in which we can find hate speech definitions (§4.1). We provide information about the respective sources and a final corpus analysis to gain cross-domain and cross-cultural insights (§4.2). Figure 1 summarizes the HateDefCon dataset creation pipeline. We collect definitions from the selected domains, we extract keywords which we use to annotate as components for each definition. HateDefCon, along with all the results of the experiments is publicly available. [^2] https://github.com/katkorre/SCA-of-Hate-Speech.git
@@ 4.1 Data
We source definitions from the Global Handbook of Hate Speech Laws website, [^3] https://futurefreespeech.org/global-handbook-on-hate-speech-laws/. a comprehensive resource that provides access to existing hate speech legislation from countries worldwide, including the United Nations and European Union levels. The legislations are already available in English. We exclude countries for which their legislative texts were not available. The full list of the countries is available in Appendix B.
@@
We use the Hate Speech Dataset Catalogue  [^5] https://hatespeechdata.com/ to navigate through research articles and datasets, focusing specifically on definitions of hate speech. We do not include other related terms like toxicity or abusive language, as our primary focus is hate speech.
@@ 4.2 Component Annotation
To ensure objectivity in the annotation process and a very fine-grained representation of the components, we establish that a definition contains a given component if a derivative of the component’s word appears in the annotation. For instance, if the word ‘abusive’ is present in the definition, the annotator marks the component ‘abuse’ as present. A challenge we encounter during annotation is the treatment of synonyms. For instance, the words ‘harm’ and ‘hurt’ can be considered synonymous in the context of hate speech definitions, but they are not derivatives of the same root word.777‘Harm’ comes from Middle English ’harm’, while ’hurt’ comes from Middle English ‘hurten’ To maintain consistency and avoid subjective interpretation, which could lead to different grouping of the synonyms by the annotators, we only annotate derivatives of the target term (e.g., marking ‘harm’ only if words like ‘harmful’ or ‘harmed’ appear). Additionally, missing components are labeled as ‘undefined component’. This process allows us to comprehensively gather all components deemed crucial by the annotators.
To finalize the dataset, we keep all annotated components, as a manual inspection showed that most of the disagreement derived from the fact that one of the annotators missed a component. The average inter-annotator agreement (IAA) measured via Cohen’s kappa McHugh (2012) is 0.64 (see Appendix D for more details). Figure 2 reports the component hierarchy resulting from the annotation study (see Appendix E for a fine-grained version of the hierarchy). Figure 3 reports the overall frequency for the top 20 components. The most frequent target components are those mostly related to racism: religion, race, ethnicity and nationality. This indicates a stronger emphasis on aspects of identity that hold deep historical and social significance in discriminatory narratives, compared to biological attributes such as age and sex, which, while still present, appear less prominently.
Figure 3: Top 20 most frequent components across all definitions.
@@ 5 Case Study: Definition Comparison
Domain No. Definitions Culture Distribution Components
@@
Table 1: Number of definitions, culture distribution, and top 5 components per domain.
@@
We employ SCA to describe potential inconsistencies between collected hate speech definitions and the cultural reality through a real case study. Mulki et al. (2019) develop a Levantine Hate Speech and Abusive Twitter dataset in an attempt to bring into the spotlight less-spoken Arabic varieties. However, in their paper, the authors refer to the hate speech definition by Nockleby (2000), which defines hate speech as “any communication that disparages a person or a group on the basis of some characteristic such as race, color, ethnicity, gender, sexual orientation, nationality, religion, or other characteristic”. Levantine Arabic is spoken in many Middle Eastern countries, such as Syria, Jordan, and Israel. Our SCA framework allows us to see if the definitions available for these countries overlap with the one used in Mulki et al. (2019) (Table 2). While they all differ in the intent and act, some targets overlap, although the definition from Mulki et al. (2019) is more comprehensive, considering also sexual orientation and gender, which are not taken into account by the other definitions. Israel’s legislation seems to have the most detailed provisions, including concepts of hostility and cultural control, whereas Syria’s legislation is narrowly focused on race, and Jordan’s legislation emphasizes physical action and terrorism. This variation can affect annotation and prompting, resulting in varied interpretations of hate-related content. Therefore, tailoring prompts to align with specific regional definitions is essential for achieving consistent model behavior for a specific culture.
@@ 6 LLM Sensitivity to Definitions
We experiment with three open-source popular state-of-the-art LLMs: Llama3888Llama3-8B-instruct., Mistral2999Mistral-7B-Instruct-v0.2., and Phi3-mini. [^10] Phi-3-mini-4k-instruct. We consider the Gab Hate Corpus (GHC) on online hate speech conversations Kennedy et al. (2022). GHC consists of 27,665 posts from the social network service gab.com and is annotated for the presence of “hate-based rhetoric” by a minimum of three annotators. The posts are annotated based on a coding typology created by synthesizing definitions of hate speech from legal precedents, existing hate speech coding frameworks, and insights from psychology and sociology. This typology includes hierarchical labels that denote dehumanizing and violent speech, as well as indicators related to targeted groups and rhetorical framing. This dataset captures various dimensions of hate speech, making it well-suited for a detailed and fine-grained analysis and testing with multiple definitions. Although GHC is annotated in a multi-class fashion, we consider a binary setting: hate vs not hate. To limit computational overhead, we select 500 instances from the corpus, equally divided between the two classes.
@@
We perform our experiments in a zero-shot setting to understand the impact of different definitions (i.e. different components) on model performance without any additional fine-tuning or prompting strategies. To prompt the models, we use three definitions of hate speech. We select the definitions deliberately after manual inspection, aiming to assess whether the varying degrees of component coverage influence the prompting outcomes. This approach includes one definition with medium component coverage, a highly detailed one, and a very general one. We use three definitions: Dghc comes from the GHC dataset, it includes several components but is not the most comprehensive one; Dwiki is the definition provided by the Macedonian Wikipedia page, it offers a more detailed and varied perspective; and Ddict is a general definition from the Merriam-Webster dictionary (see Appendix F for more details). We set the temperature to zero to exclude variations in the generated responses.
@@ 7 Results
Table 3 summarizes the results of the classification task according to the three definitions. Although Llama3 performs consistently better than other models, the complexity of the definitions affects the outputs of the three models differently. In Llama3, the comprehensiveness of the definition is directly proportional to the performance: the more comprehensive the definition is, the more the performance increases; in Mistral2, we observe a reverse tendency. Phi3-mini outperforms Mistral2 in all settings, although being half the size of Mistral2. Additionally, in Phi3-mini, we observe the smallest variety of responses based on the definitions, with Dghc performing the best.
There are cases in which Llama3 refuses to answer for safeguarding the generation process over harmful content. This tendency is also proportional to the completeness of the definition. The less comprehensive the definition, the more the model refuses to answer. For Dwiki, the model refuses to answer to 1% of times, 1.6% for Dghc, and 5% for Ddict.
@@
Table 3: Classification performance (F1score) using difference hate speech definitions.
@@
We extract all instances for which the answer of Llama3 varies according to the definition, regardless of the ground truth. Dghc and Dwiki differ in 54 instances, while Ddict differs from Dghc and Dwiki in 48 instances each. We observe that when the model is prompted with Dwiki, it tends to identify more personal attacks, which are often discarded in Dghc and Ddict as false negatives since such definitions identify group of people as targets of hate speech, rather than individuals. Moreover, with Dwiki, LLMs tend to identify more instances that are hateful with respect to political views and undocumented migrants. Indeed, the majority of instances that are correctly identified by Dwiki and misclassified by Dghc and Ddict, contain terms like liberals, communists, and illegal aliens with a negative connotation. In short, we see that the models indeed carry their own biases which are evident in how they classify instances even if we change the definition in the prompt, indicating that they rely more on their own internal knowledge than solely on the definition.
@@ 8 Conclusions
Acknowledgments
K. Korre’s research is carried out under the project RACHS: Rilevazione e Analisi Computazionale dell’Hate Speech in rete, in the framework of the PON programme FSE REACT-EU, Ref. DOT1303118. F. Ruggeri is partially supported by the project European Commission’s NextGeneration EU programme, PNRR – M4C2 – Investimento 1.3, Partenariato Esteso, PE00000013 - “FAIR - Future Artificial Intelligence Research” – Spoke 8 “Pervasive AI”. Arianna Muti’s research is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101116095, PERSONAE).